# Revisiting 360 Depth Estimation with PanoGabor: A New Fusion Perspective

### v1

@@ Abstract
Depth estimation from a monocular 360 image is important to the perception of the entire 3D environment. However, the inherent distortion and large field of view (FoV) in 360 images pose great challenges for this task. To this end, existing mainstream solutions typically introduce additional perspective-based 360 representations (e.g., Cubemap) to achieve effective feature extraction. Nevertheless, regardless of the introduced representations, they eventually need to be unified into the equirectangular projection (ERP) format for the subsequent depth estimation, which inevitably reintroduces the troublesome distortions. In this work, we propose an oriented distortion-aware Gabor Fusion framework (PGFuse) to address the above challenges. First, we introduce Gabor filters that analyze texture in the frequency domain, thereby extending the receptive fields and enhancing depth cues. To address the reintroduced distortions, we design a linear latitude-aware distortion representation method to generate customized, distortion-aware Gabor filters (PanoGabor filters). Furthermore, we design a channel-wise and spatial-wise unidirectional fusion module (CS-UFM) that integrates the proposed PanoGabor filters to unify other representations into the ERP format, delivering effective and distortion-free features. Considering the orientation sensitivity of the Gabor transform, we introduce a spherical gradient constraint to stabilize this sensitivity. Experimental results on three popular indoor 360 benchmarks demonstrate the superiority of the proposed PGFuse to existing state-of-the-art solutions. Code can be available upon acceptance.
@@
Monocular, 360°, Depth Estimation, Distortion, Gabor Transform
Figure 1: Comparative of our proposed framework (shown in red) against other existing methods. Our method consistently demonstrates lower values across all error metrics. The compact red area highlights that our approach maintains competitive scores in computational efficiency (FLOPs), demonstrating a well-balanced trade-off between performance and computational cost.
@@ I Introduction
To mitigate the negative effects of distortion, BiFuse [1] introduced a dual-projection fusion scheme, which captures distortion-free features using an additional cubemap representation. However, its bidirectional fusion module imposes significant computational and memory burdens. To address this issue, Unifuse [2] proposes a more efficient unidirectional fusion module based on residual learning to reduce computational demands. Moreover, HRDfuse [3] replaces the cubemap projection with a more advanced Tangent Patch projection format (TP), achieving an effective fusion of TP and ERP features through a spatial feature alignment module. On the other hand, some researchers focus on introducing distortion awareness to transformer architectures with large receptive fields [4, 5, 6]. Considering the non-negligible computational and memory cost inherent in transformers, Elite360D [7] proposes a bi-projection bi-attention fusion module that integrates ERP and icosahedron projection (ICOSAP) features, combining the strengths of fusion-based schemes and transformer-based architectures. However, this type of fusion-based method, whether introducing new representations or proposing new fusion modules, overlooks the fact that all representations need to be integrated into the ERP format for fusion, which inevitably reintroduces distortion. Therefore, we argue that explicitly addressing the troublesome panoramic distortion in the fusion process can benefit the subsequent depth estimation.
@@
To this end, we present PGFuse, an oriented panoramic Gabor-based fusion framework for monocular omnidirectional depth estimation. PGFuse addresses reintroduced distortions when integrating other representations into the ERP format. Benefiting from the frequency domain transformation, Gabor filters can extend the receptive field. Meanwhile, due to the strength of texture analysis capability, it can reduce the interference of surface texture, thereby enhancing the geometric depth cues. Besides, the tunable parameters of Gabor filters enable the creation of task-specific filters without additional parameters. However, the inherent distortion in 360 images poses a significant challenge for Gabor filter applications. To address this challenge, we design distortion-aware Gabor filters using a linear latitude-aware distortion representation method, as detailed in Eq. 2 and Algorithm 1, which adjusts the parameters of Gabor filters to achieve distortion awareness. These finely tuned Gabor filters, termed PanoGabor filters, are then incorporated as a convolutional layer. PanoGabor is embedded into a specially designed Channel-wise and Spatial-wise Unidirectional Fusion Module (CS-UFM) to mitigate the reintroduced distortions. More importantly, the parameters of PanoGabor can be optimized via dataset-driven training, ensuring seamless integration. Furthermore, we introduce a spherical gradient technique to stabilize the orientation sensitivity of the Gabor filters. Leveraging Gabor-based distortion-aware design and a perspective-based feature extraction strategy, PGfuse achieves the best-balanced trade-off between performance and computational cost, as shown in Figure. 1.
@@
Extensive experiments evaluated on three popular indoor datasets—Stanford2D-3D [8], Matterport3D [9], and Structured3D [10]—demonstrate that our proposed scheme outperforms other state-of-the-art approaches. Our contributions can be summarized as follows.
@@
  • We design a channel-wise and spatial-wise unidirectional fusion module for the integration of other representations into ERP format, delivering effective and distortion-free features for the subsequent depth map estimation.
  • In popular benchmarks, our solution outperforms current state-of-the-art approaches.
Figure 2: Overview of the proposed framework.
@@ II-A Monocular 360 Depth Estimation
Panoramic distortion leads to notable performance degradation in perspective image-based depth estimation approaches. To this end, Zioulis et al. [11] highlighted the need to learn within the 360 domain, introducing the 3D60 dataset. Moreover, to mitigate distortion, researchers have adjusted the sampling positions of convolutional kernels for distortion awareness [12, 13, 14]. Some indoor monocular panoramic depth estimation methods [15, 16, 17] have developed various distortion-aware techniques to manage panoramic distortions, but most of these techniques are computational inefficient. Recently, advanced Transformer-based architectures [18] have broadened the receptive field, enabling models to extract features from 360 images. However, the traditional self-attention mechanism struggles to handle the complex distortion distribution in 360 images [4]. Enhanced distortion-aware transformer architectures [4, 5, 6, 19], while effective, often come with high computational and memory costs. To simplify models, some approaches [20, 21] advocate for estimating indoor depth from vertically compressed one-dimensional sequences, which leverages gravity alignment properties. However, this compression often fails to capture essential structural geometric information. To achieve a balanced trade-off between performance and efficiency, the mainstream approaches still prefer to develop a more efficient fusion-based framework. However, these fusion methods neglect the reintroduced distortion during the fusion process. Besides, although many of these approaches claim state-of-the-art performance, comparisons are often complicated by variations in datasets, preprocessing methods, and evaluation strategies, making fair comparisons challenging. To this end, Elite360D [7] provides a solid baseline. Therefore, in this paper, we follow the Elite360D method [7] to conduct the experiments.
II-B Gabor Filter in Image Processing
The Gabor filter is widely used in fields such as texture analysis, object detection, and face recognition, due to its robustness to varying angles and scales. Li et al. [22] suggested enhancing face recognition by learning the Gabor wavelet’s covariance matrix. In fabric defect detection, Chen et al. [23] introduced the Genetic Algorithm Gabor Faster R-CNN (Faster GG R-CNN), embedding Gabor kernels to reduce texture interference. Andrey et al. [24] proposed GaborNet, which uses Gabor filters to improve convergence and reduce the complexity of training in image recognition. To extend the receptive fields and enhance the geometric depth cues, we introduce Gabor filters to analyze texture in the frequency domain. Considering the special characteristics of 360 images, we propose a distortion-aware Gabor-based approach to address the challenging distortion issues.
@@ III Methodology
In this section, we introduce the proposed PGFuse network. The complete framework, as illustrated in Figure 2, takes an ERP image as input and produces the corresponding predicted depth map. Our model follows a typical encoder-decoder scheme and consists of three main components: a feature extractor, a unidirectional fusion module, and an up-sampling decoder. To reduce the computational burden, we employ a shared-weight backbone to extract two sets of distortion-free cubemap features: one from the original perspective and the other horizontally rotated by 45°. These features are spatially aligned, unified into the ERP format, and then inputted into the CS-UMF module.
@@
In the fusion module, we design PanoGabor filters to effectively handle the reintroduced distortion and enhance depth cues by analyzing textures in the frequency domain. Specifically, we introduce a linear latitude-aware distortion representation method that approximates the cosine value of latitude with a linear projection strategy. We then design an Algorithm 1 that uses this distortion representation to generate distortion-aware PanoGabor filters. These filters are further enclosed as a convolution layer and embedded in the CS-UFM module. Subsequently, the features processed by the CS-UFM module are fed into a multi-scale distortion-aware module (MS-DAM) [25] to integrate the shallow geometric features and the deep semantic features. These integrated features are then fed into the decoder, following a hierarchical architecture with multi-scale stages. Each stage consists of an upsampling layer and a convolution layer, followed by an ELU activation function. A final 3×3 convolution layer is used to produce the final depth map. Additionally, we project the ERP depth map onto a spherical surface and employ the Sobel operator to compute depth gradients in the nearest tangent patch, providing additional constraints and stabilizing the orientation sensitivity of the PanoGabor filters. The following sections provide detailed descriptions of the proposed modules.
@@
Input: $c$: The defined distortion coefficient
Output: Convolution parameters
1:  Initialize convolution layer with given parameters
@@
2:  Initialize Gabor parameters:$\theta\leftarrow\{\frac{\pi\cdot i}{8}\}_{i=1}^{8}$; $\psi=\pi\epsilon$; $f\leftarrow(\frac{\pi}{2})\cdot(\sqrt{2})^{\epsilon}\cdot(1+c)$; $\sigma\leftarrow\frac{\pi}{f+0.1}$
3:  Create meshgrid for $y$ and $x$: $[y,x]\leftarrow\text{meshgrid}((-1,0,1),(-1,0,1))$
4:  Reshape $\theta$, $\sigma$, $f$, and $\psi$ to match the required shape
5:  Calculate rotated $x$ and $y$: $x_{rot}\leftarrow x\cdot\cos(\theta)+y\cdot\sin(\theta)$; $y_{rot}\leftarrow-x\cdot\sin(\theta)+y\cdot\cos(\theta)$
@@
6:  Calculate Gabor filter $g$:$g\leftarrow\exp\left(-0.5\cdot\left(\frac{x_{rot}^{2}+y_{rot}^{2}}{\sigma^{2}}\right)\right)\cdot\cos(f\cdot x_{rot}+\psi)$$g\leftarrow\frac{g}{2\cdot\pi\cdot\sigma^{2}}$
7:  Set convolutional parameters $\leftarrow g$
8:  return  Initialized distortion-aware Gabor-based convolutional parameters
III-A Distortion-aware PanoGabor Filters (PGF)

### v4

@@ Abstract
Depth estimation from a monocular 360 image is important to the perception of the entire 3D environment. However, the inherent distortion and large field of view (FoV) in 360 images pose great challenges for this task. To this end, existing mainstream solutions typically introduce additional perspective-based 360 representations (e.g., Cubemap) to achieve effective feature extraction. Nevertheless, regardless of the introduced representations, they eventually need to be unified into the equirectangular projection (ERP) format for the subsequent depth estimation, which inevitably reintroduces additional distortions. In this work, we propose an oriented-distortion-aware Gabor Fusion framework (PGFuse) to address the above challenges. First, we introduce Gabor filters that analyze texture in the frequency domain, extending the receptive fields and enhancing depth cues. To address the reintroduced distortions, we design a latitude-aware distortion representation to generate customized, distortion-aware Gabor filters (PanoGabor filters). Furthermore, we design a channel-wise and spatial-wise unidirectional fusion module (CS-UFM) that integrates the proposed PanoGabor filters to unify other representations into the ERP format, delivering effective and distortion-aware features. Considering the orientation sensitivity of the Gabor transform, we further introduce a spherical gradient constraint to stabilize this sensitivity. Experimental results on three popular indoor 360 benchmarks demonstrate the superiority of the proposed PGFuse to existing state-of-the-art solutions. Code and models will be available at https://github.com/zhijieshen-bjtu/PGFuse.
@@
Monocular, 360°, Depth Estimation, Distortion, Gabor Transform.
Figure 1: Comparison of our proposed framework (shown in red) against other existing methods. Our method consistently demonstrates lower values across all error metrics. The compact red area highlights that our approach maintains competitive scores in computational efficiency (FLOPs), demonstrating a well-balanced trade-off between performance and computational cost.
@@ I Introduction
To mitigate the negative effects of distortion, BiFuse [1] introduced a dual-projection fusion scheme, which captures distortion-aware features using an additional cubemap representation. However, its bidirectional fusion module imposes significant computational and memory burdens. To address this issue, Unifuse [2] proposes a more efficient unidirectional fusion module based on residual learning to reduce computational demands. Moreover, HRDfuse [3] replaces the cubemap projection with a more advanced Tangent Patch projection format (TP), achieving an effective fusion of TP and ERP features through a spatial feature alignment module. On the other hand, some researchers focus on introducing distortion awareness to transformer architectures with large receptive fields [4, 5, 6]. Considering the non-negligible computational and memory cost inherent in transformers, Elite360D [7] proposes a bi-projection bi-attention fusion module that integrates ERP and icosahedron projection (ICOSAP) [8, 9] features, combining the strengths of fusion-based schemes and transformer-based architectures. However, this type of fusion-based method, whether introducing new representations or proposing new fusion modules, overlooks the fact that all representations need to be integrated into the ERP format for fusion, which inevitably reintroduces distortion. Therefore, we argue that explicitly addressing the troublesome panoramic distortion in the fusion process can benefit the subsequent depth estimation.
@@
To this end, we present an oriented-distortion-aware Gabor Fusion framework (termed PGFuse), an oriented panoramic Gabor-based fusion framework for monocular omnidirectional depth estimation. PGFuse can address reintroduced distortions when integrating other representations into the ERP format. Specifically, we explore the Gabor wavelets for panoramic depth estimation, aiming to leverage the advantages of a human vision-inspired model. Recent research has focused on extracting effective features based on human visual principles [10, 11]. The Gabor wavelet kernel, resembling the two-dimensional receptive field of human vision, is highly effective at capturing key visual features such as spatial positioning, directional selectivity, and spatial frequency. Unlike other methods that extend the receptive field, Gabor filters are particularly effective in panoramic depth estimation due to their ability to capture multi-scale spatial frequencies and directional information simultaneously. This enables them to handle better the large, varying depth ranges and complex textures typically encountered in panoramic scenes, improving depth estimation accuracy. Besides, the tunable parameters of Gabor filters enable the creation of task-specific filters without additional parameters. However, the inherent distortion in 360 images poses a significant challenge for Gabor filter applications. To address this challenge, we design Gabor filters using a latitude-aware distortion representation method, which adjusts the parameters of Gabor filters to achieve distortion awareness. These finely tuned Gabor filters, termed PanoGabor filters, are then incorporated as a convolutional layer. PanoGabor is embedded into a specially designed Channel-wise and Spatial-wise Unidirectional Fusion Module (CS-UFM) to mitigate the reintroduced distortions. More importantly, the parameters of PanoGabor can be optimized via dataset-driven training, ensuring seamless integration. Furthermore, we introduce a spherical gradient technique to stabilize the orientation sensitivity of the Gabor filters. Leveraging Gabor-based distortion-aware design and a perspective-based feature extraction strategy, PGfuse achieves the best-balanced trade-off between performance and computational cost, as shown in Fig. 1.
@@
Extensive experiments evaluated on three popular indoor datasets Stanford2D-3D [12], Matterport3D [13], and Structured3D [14] demonstrate that our proposed scheme outperforms the state-of-the-art approaches. Our contributions can be summarized as follows:
@@
  • We design a channel-wise and spatial-wise unidirectional fusion module for the integration of other representations into ERP format, delivering effective and distortion-aware features for the subsequent depth map estimation.
Figure 2: Overview of the proposed framework. Our framework takes a single panorama as input and outputs the corresponding depth map. This network addresses the reintroduced distortion problems that are often overlooked by previous fusion-based methods. Specifically, we introduce a PanoGabor approach to achieve the distortion awareness.
@@ II-A Monocular 360 Depth Estimation
Panoramic distortion leads to notable performance degradation in perspective image-based depth estimation approaches. To this end, Zioulis $et~{}al.$ [15] highlighted the need to learn within the 360 domain, introducing the 3D60 dataset. Moreover, to mitigate distortion, researchers have adjusted the sampling positions of convolutional kernels for distortion awareness [16, 17, 18]. Some indoor monocular panoramic depth estimation methods [19, 20, 21] have developed various distortion-aware techniques to manage panoramic distortions, but most of these techniques are computationally inefficient. Recently, advanced Transformer-based architectures [22] have broadened the receptive field, enabling models to extract features from 360 images. However, the traditional self-attention mechanism struggles to handle the complex distortion distribution in 360 images [4]. Enhanced distortion-aware transformer architectures [4, 5, 6, 23, 24, 25], while effective, often come with high computational and memory costs. To simplify models, some approaches [26, 27, 28, 29, 30] advocate for estimating indoor depth from vertically compressed one-dimensional sequences, which leverages gravity alignment properties. However, this compression often fails to capture essential structural geometric information. To achieve a balanced trade-off between performance and efficiency, the mainstream approaches[31, 32, 7] still prefer to develop a more efficient fusion-based framework. However, these fusion methods neglect the reintroduced distortion during the fusion process. Besides, although many of these approaches claim state-of-the-art performance, comparisons are often complicated by variations in datasets, preprocessing methods, and evaluation strategies, making fair comparisons challenging. To this end, Elite360D [7] provides a solid baseline. Therefore, in this paper, we follow the Elite360D method [7] to conduct the experiments.
II-B Gabor Filters in Image Processing
The Gabor filter is widely used in texture analysis, object detection, and face recognition due to its robustness to varying angles and scales. Li $et~{}al.$ [33] suggested enhancing face recognition by learning the Gabor wavelet’s covariance matrix. In fabric defect detection, Chen $et~{}al.$ [34] introduced the Genetic Algorithm Gabor Faster R-CNN (Faster GG R-CNN), embedding Gabor kernels to reduce texture interference. Andrey $et~{}al.$ [35] proposed GaborNet, which uses Gabor filters to improve convergence and reduce the complexity of training in image recognition. To extend the receptive fields and enhance the geometric depth cues, we introduce Gabor filters to analyze texture in the frequency domain. Considering the special characteristics of 360 images, we propose a distortion-aware Gabor-based approach to address the challenging distortion issues.
@@ III Methodology
In this section, we introduce PGFuse, its pipeline is illustrated in Fig. 2. It takes an ERP image as input and produces the corresponding predicted depth map. Our model follows a typical encoder-decoder scheme and consists of three main components: a feature extractor, a unidirectional fusion module, and an up-sampling decoder. To reduce the computational burden, we employ a shared weight backbone to extract two sets of distortion-free cubemap features: one from the original perspective and the other horizontally rotated by 45°. The two groups of cube features are employed to relieve the pixel-missing and discontinuous boundary problems inherent in the cubemap projection[29]. These features are spatially aligned, unified into the ERP format, and then inputted into the CS-UFM module.
@@
In the fusion module, we design PanoGabor filters to effectively handle the reintroduced distortion and enhance depth cues by analyzing textures in the frequency domain. Specifically, we introduce a latitude-aware distortion representation method that approximates the cosine value of latitude. We then design a Distortion-Aware PanoGabor Filters Algorithm 1 that uses this approach to generate PanoGabor filters. These filters are further enclosed as a convolution layer and embedded in the CS-UFM module. Subsequently, the features processed by the CS-UFM module are fed into a multi-scale distortion-aware module (MS-DAM) [25] to integrate the shallow geometric features and the deep semantic features. These integrated features are then fed into the decoder, following a hierarchical architecture with multi-scale stages. Each stage consists of an upsampling layer and a convolution layer, followed by an ELU activation function. Finally, a 3×3 convolution layer is used to produce the final depth map. In addition, we project the ERP depth map onto a spherical surface and employ the Sobel operator to compute depth gradients in the nearest tangent patch, providing additional constraints and stabilizing the orientation sensitivity of the PanoGabor filters. The following sections provide detailed descriptions of the proposed modules.
Figure 3: Illustration of the differential area element on the spherical surface (left) and the ERP domain (right).
III-A Preliminary
Distortion Distribution Pattern. In the context of ERP projections, the distortion distribution increases progressively from the equatorial region towards the poles. To mathematically describe this distribution, we introduce a method based on the ratio of differential area elements before and after the ERP projection operation. Given a small region on the surface of a sphere, with spherical coordinates defined by the center of the sphere as the origin, we denote the differentials of latitude $\phi$ and longitude $\lambda$ by $d\phi$ and $d\lambda$, respectively. When the latitude changes from $\phi$ to $\phi+d\phi$, the corresponding arc length is $Rd\phi$, where $R$ is the radius of the sphere. At latitude $\phi$, the arc length in the longitude direction is $R\cos(\phi)d\lambda$, since the distance corresponding to a unit change in longitude decreases as latitude increases. Consequently, the differential area element of the region on the spherical surface (Fig. 3(left)) can be expressed as:
$dA=R^{2}\cos(\phi)\,d\phi\,d\lambda,$ (1)
where $\cos(\phi)$ accounts for the latitude-dependent scaling factor. However, in the ERP domain, the variations in both latitude and longitude are assumed to be uniform due to the stretching effect. As a result, the differential area element in this domain (Fig. 3(right)) is approximated as:
$dA^{\prime}=Rd\phi\,Rd\lambda=R^{2}d\phi\,d\lambda.$ (2)
The distortion factor $K(\phi)$ is then defined as the ratio of the differential area element in the ERP projection to the differential area element on the spherical surface:
$K(\phi)=\frac{dA^{\prime}}{dA}=\frac{1}{\cos(\phi)}.$ (3)
The distortion factor $K(\phi)$ ranges from $1$ to +$\infty$+ ∞. At the equator, the projection corresponds to an identity mapping, implying that the distortion is zero at this latitude. To more accurately describe the variation in distortion across latitudes, we introduce a bias term, leading to the following expression for the distortion factor:
$K(\phi)=\frac{1}{\cos(\phi)}-1.$ (4)
This formulation characterizes the distortion as a function of latitude in the ERP projection more precisely. It provides a foundation for further distortion analysis and potential compensation techniques in the context of ERP projections.
Gabor Transform. When applied to image processing, the Gabor transform can provide useful spatial frequencies and orientations analysis of the images. By responding to specific spatial frequencies and orientations, the Gabor filters can effectively identify textures and detect edges, which are critical for depth estimation. The standard Gabor transform formula can be denoted as:
$g(x,y,f,\theta,\psi,\sigma)=\frac{1}{2\pi\sigma^{2}}e^{-\frac{x^{{}^{\prime}2}+y^{{}^{\prime}2}}{2\sigma^{2}}}\cos(fx^{{}^{\prime}}+\psi),$ (5)
where
$x^{{}^{\prime}}$ $=$ $x\cos\theta+y\sin\theta$ (6)
$y^{{}^{\prime}}$ $=$ $-x\sin\theta+y\cos\theta,$ (7)
and $(x,y)$ represents the coordinates of the Gabor kernels (the pixel position in the spatial domain); the kernel size is set to 3 (i.e., $x,y\in\{-1,0,1\}$); $f$ indicates the central angular frequency of a sinusoidal plane wave, $\theta$ represents the orientation of the Gabor filter, $\psi$ is the phase offset, $\sigma$ represents the sharpness of the Gaussian function along both $x$ and $y$ directions.
@@ III-A Preliminary
Input:
$C_{in}$
: input channel numbers;
$C_{out}$: output channel numbers;
$K(\phi_{C})\in\mathbb{R}^{C_{in}}$: distortion coefficients;
Output: PanoGabor Filters
1:  Define 8 orientations: $\{\theta_{i}\in\mathbb{R}^{\frac{C_{in}}{8}\times C_{out}}=\frac{\pi\cdot i}{8}\}_{i=0}^{7}$
2:  Divide $C_{in}$ into 8 groups:
3:  for $i^{th}$ group do
4:     apply $\theta_{i}$
5:  end for
6:  $\theta\in\mathbb{R}^{C_{in}\times C_{out}}\leftarrow$ concatenate
@@
7:  Calculate Gabor parameters [^1] The parameters are set up according to [11], $\epsilon$ is generated randomly.:$\psi\in\mathbb{R}^{C_{in}\times C_{out}}\leftarrow\pi\epsilon$; $f\in\mathbb{R}^{C_{in}\times C_{out}}\leftarrow(\frac{\pi}{2})\cdot(\sqrt{2})^{\epsilon}$;$\sigma\in\mathbb{R}^{C_{in}\times C_{out}}\leftarrow\frac{\pi}{f+0.1}$
8:  $f\leftarrow f\cdot(1+K(\phi_{C}))$;
9:  $(y,x)\leftarrow$ meshgrid((1,0,1),(1,0,1))101101((-1,0,1),(-1,0,1))( ( - 1 , 0 , 1 ) , ( - 1 , 0 , 1 ) ); $x,y\in\mathbb{R}^{3\times 3}$
10:  Rotate $x$ and $y$:$x_{rot}\leftarrow x\cdot\cos(\theta)+y\cdot\sin(\theta)$;$y_{rot}\leftarrow-x\cdot\sin(\theta)+y\cdot\cos(\theta)$
@@
11:  Calculate Gabor filter $g$:$g\leftarrow\exp\left(-0.5\cdot\left(\frac{x_{rot}^{2}+y_{rot}^{2}}{\sigma^{2}}\right)\right)\cdot\cos(f\cdot x_{rot}+\psi)$;$g\leftarrow\frac{g}{2\cdot\pi\cdot\sigma^{2}}$
12:  Reshape $g\in\mathbb{R}^{C_{in}\times C_{out}\times 3\times 3}$
13:  return  $g$
III-B Distortion-aware PanoGabor Filters (PGF)